OpenStudy (anonymous):

write an equation of the lien containing th given point and parallel to the given line (-4,4) 9x=4y+3

OpenStudy (anonymous):

parallel lines have the same slope. lets rewrite the given equation in the slope intercept form. do that and post what you get.

OpenStudy (anonymous):

Parallel lines have the same slope, so the slope of the parallel line can be determined from the equation you are given. slope = 9/4 y = (9/4)x + instercept substitute the point (-4,4) in this equation (y = (9/4)x + instercept) to determine the value of the intercept. slope = \cfrac{9}{4} 4 = \left(\cfrac{9}{4}\right)\left(-4\right)+intercept The solution is attached.

OpenStudy (anonymous):

Oops... getting the hang of the type setting...I meant: \[slope = \cfrac{9}{4}\] \[4 = \left(\cfrac{9}{4}\right)\left(-4\right)+intercept\]

OpenStudy (anonymous):

Parallel lines have the same coefficients of x and y, constant term is different. The equation of the line containing the given point and parallel to the given line (-4,4) 9x=4y+3 will be 9x = 4y +c Plug in x =-4, y =4 and solve for c 9*(-4) = 4*4 +c -36 = 16 + c -52 =c Equation of line is 9x = 4y -52